OpenStudy (anonymous):

Add: - (5 )/6 + (2 )/3. Express the result as a fraction.

OpenStudy (mathstudent55):

What is the least common denominator of 3 and 6?

OpenStudy (anonymous):

a worthy question, 100/12+ 4/12.

OpenStudy (mathstudent55):

I'd say it's 6. -5/6 + 2/3 = -5/6 + 2/3*2/2 = -5/6 + 4/6 = -1/6

OpenStudy (anonymous):

uh huh.

OpenStudy (anonymous):

:) good dat, mathstudent55

OpenStudy (anonymous):

ew, day.

OpenStudy (anonymous):

Does it mean something when the number is in parentheses like multiplying?

OpenStudy (mathstudent55):

In this case, it doesn't.